OpenStudy (anonymous):

the three math idols judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 610! If four times the number of one-step equations minus twice the number of two-step equations is equal to 350, how many one-step equations auditioned today? A-260 B-350 C-340 D-670

OpenStudy (anonymous):

one+two=610 4*one-2*two=350

OpenStudy (anonymous):

one=610-two sfone plug it into 2nd eq 4*(610-two)-2*two=350 solve for two

OpenStudy (anonymous):

please please tell me where this problem comes from. no one will say, but i see it more than once a week.

OpenStudy (anonymous):

two is 1045/3

OpenStudy (anonymous):

plug it back into first equation 1045/3+one=610 solve for one

OpenStudy (anonymous):

why you want to know though?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

i get 261.67 but round down to 260

OpenStudy (anonymous):

I think he is just curious